OpenStudy (azureflame241):

How do you simplify x-3/(x^4-81) ?

OpenStudy (anonymous):

\[\frac{ x-3 }{ (x^2+9)(x^2-9) }=\frac{ x-3 }{ (x^2+9)(x+3)(x-3) }=\frac{ 1}{ (x^2+9)(x+3) }\]

OpenStudy (azureflame241):

Thank you!